OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (0, 0) m = 2 I need help understanding how to set up the equation.

OpenStudy (anonymous):

The equation you want should look like this: \(y=mx+b\). You're told that \(m=2\), so \(y=2x+b\). You know the line passes through the origin, (0,0). This means when \(x=0\) you have \(y=0\). Plug this into the equation above: \[0=2(0)+b~~\Rightarrow~~b=0\] So, the equation of the line is \(y=2x\).

OpenStudy (anonymous):

Does that formula work for any variables? ex. (5,5) m = -3 ?

OpenStudy (anonymous):

Yes, the equation of a line depends on the slope and any known point on the line. Given (5,5) and \(m=-3\) (the point and slope respectively), you would have \[5=-3(5)+b~~\Rightarrow~~b=20\]and\[y=-3x+20 \]

OpenStudy (anonymous):

Okay thank you very much!

OpenStudy (anonymous):

You're welcome